OpenStudy (anonymous):

Which of the following are true statements? (Check all that apply) a. logM+logN=log(MN) b. logM^p=P*logM c. (logM)^p=P*logM d. logM-logN=log(M-N)

OpenStudy (anonymous):

first is true for sure, one of the basic properties of the log

OpenStudy (anonymous):

second is also true, assuming it means \[\log(M^p)=p\log(M)\]

OpenStudy (anonymous):

last two are waaaay wrong

OpenStudy (anonymous):

Yes thats what it says, couldnt really put as they were on here tho

OpenStudy (anonymous):

Okay Thank you!